# ppa_t_p.bib

### v1

@@ 4 Anonymous Credentials: The Theory
We now first want to clarify what we mean by “privacy-preserving authentication”. Essentially, this is a signature-based authentication approach that supports $i)$ selective disclosure and $ii)$ unlinkability in all the flavors that we have discussed. Consequently, even if IdP and RPs collude, it is not be possible to track users. Note that this makes the user decide which information (e.g., attributes) are revealed and provide a strong unlinkability guarantee by default. As already mentioned in Section 1, there are different types of anonymous credentials with a set of different features as well as (closely) related primitives (cf. [DBLP:conf/secsr/KakviMPQ23, cryptoeprint:2023/1039]). In the following we will consider a simple abstract concept and will then discuss different (basic) features in Section 4.2.
@@ 4.1 The Abstract Concept
IssuerUserVerifier 1Verifier 2Issuing (attributes $A$)Showing of $A_{1}\subseteq A$Showing of $A_{2}\subseteq A$///Showings are unlinkable
@@
At the end of this interaction, the user has a credential, which can be seen as the signature of the issuer on attributes $A$. Then, the user can perform a showing protocol with verifiers, where in each interaction a subset $A_{i}\subseteq A$ of the attributes in the credential can be revealed. For simplicity, here we use a simple selective disclosure of attributes, but this could actually be a more complex predicate over a subset of attributes (as discussed in Section 4.2). For instance, a user might want to consume some content online where they need to demonstrate that they satisfy some age rating or that they are residents of some country. Note that in those cases revealing the full identity of the user is typically not necessary. If a verifier accepts such a showing protocol, then it can be sure it interacts with a user that has a valid credential from the issuer for the subset of attributes that were revealed during the showing protocol. Note that as long as none of the attributes shown are identifying and there is more than one user who could potentially also reveal the same set of attributes, the user stays anonymous. Note that in Figure 2 we can also have that both verifiers are the same entity.
@@
Security. What we request from an anonymous credential scheme is that it is unforgeable, that is, it is not possible for a user to demonstrate some set of attributes for which they did not obtain a credential. Moreover, the scheme is unlinkable, which means that it is not possible to link together the showings of the same user. We note that for any candidate scheme considered to be deployed in practice it is imperative to come with a rigorous proof of security in a realistic security model.
@@
Extensions to this simple model. As briefly mentioned above, in realistic scenarios usually there will be more than one issuer. Such systems are sometimes called multi-authority anonymous credential systems. In such a setting, every issuer has an independent signing key to issue credentials and a showing might then involve more than one credential. Such an extension is typically straightforward, but if one targets at specific properties, e.g., constant credential showings irrespective of the number of issuers, dedicated constructions exist [DBLP:conf/scn/HebantP22, DBLP:conf/ccs/MirBGLS23]. We note that today there is a trend to move away from this centralized setting towards a decentralized identity. A popular concept in this decentralized identity space is that of self-sovereign identity (SSI) with Sovrin [^5] https://sovrin.org being a prominent example. Moreover, the European Union Digital Identity Wallet (EUDIW) later discussed in Section 6 can be viewed as an approach towards SSI. In SSI users are collecting certified attributes (called verifiable credentials) from different sources and then presenting (subsets of) verifiable credentials from this collection.
@@
Another flavor of having multiple issuers is that of threshold issuance anonymous credentials [DBLP:conf/ndss/SonninoABMD19]. Here, rather than every issuer holding an independent signing key, all the credentials are issued under a single signing key. However, the key is shared among all the issuers using a $(k,n)$-threshold secret sharing. In such a setting there are $n$ issuers each holding a share of the secret key and $k$ issuers need to collaborate to be able to issue a credential. Note that this is a means to distribute the trust in the issuers and adds security, as it is required that even leaking $t<k$ secret key shares does not allow to learn the signing key. Moreover, it adds robustness as the availability of any $k$ out of the $n$ issuers can issue credentials.
@@ 4.2 Features of Anonymous Credentials
In the following we want to present some basic and distinguishing features of anonymous credential systems and also provide a list of extended features. We try to be comprehensive in our presentation, but do not claim to be exhaustive. See also [DBLP:journals/istr/CamenischDELNPP14] for a discussion of features.
@@
In the single-use case if a credential is used more than once, different showings become linkable. Multi-use credentials support an unlimited number of unlinkable showings. When talking about anonymous credentials, one typically implicitly assumes multi-use ones.
@@
Typically it should be possible to encode attributes into an anonymous credential. However, over the recent years some use-cases appeared in industry in which attributes are not really required. The latter concept is then typically called anonymous tokens starting with Privacy Pass [DBLP:journals/popets/DavidsonGSTV18], which are build from blind signatures or related primitives such as oblivious pseudo-random functions (OPRFs).
@@
An issue with any digital credential is that it can easily be copied and thus illegally distributed. This gets particularly delicate in case of privacy-preserving authentication mechanisms where the identity will not be revealed. While there is no panacea, a typical approach is to bind it to hardware [DBLP:conf/crypto/Brands93, DBLP:conf/ccs/HanzlikS21], as done within the related primitive of direct anonymous attestation (DAA) [DBLP:conf/ccs/BrickellCC04] built within every Trusted Platform Module (TPM), available in most modern computers nowadays. An alternative approach is to force users to share a valuable secret whenever they share their credentials [DBLP:conf/eurocrypt/CamenischL01], a concept which however seems rather hard to implement in practice. We stress that even when credentials are bound to hardware and thus a device, this does not yet prevent the sharing of the device, e.g., lending it to a family member or friend. There are recent approaches to involve biometrics to prevent this problem. Hesse et al. suggest to use a dedicated hardware token that displays a photo of the credential holder and which is bound to an anonymous credential [DBLP:conf/uss/HesseSS23]. Adams suggested to directly integrate a biometric template as attribute into the anonymous credential [DBLP:journals/scn/Adams11]. While this approach has recently been shown to be practical [DBLP:journals/compsec/RodriguezKS24], it is best suited for physical access control (or requires additional trust assumptions).
Expressiveness of attribute showings.
Credentials may either only allow to reveal or withhold attributes (this can be thought of redacting the non-shown attributes) or be able to prove arbitrary statements about attributes encoded in the credential (e.g., attribute country represents any of the EU countries and attribute birthdate is so that that the holder is above 18 years old). Latter requires the use of zero-knowledge proofs (cf. Section 4.3), while from a construction perspective former can be constructed without them.
@@
Typically, credentials are publicly verifiable using the public key of the issuer. However, there are also constructions [DBLP:conf/ccs/ChaseMZ14] and use-cases, e.g., private groups in the Signal messenger [DBLP:conf/ccs/ChasePZ20], where the issuer is identical to the verifier and verification uses the secret of the issuer. This can allow for more efficient constructions that can avoid the use of bilinear groups and thus specific pairing-friendly elliptic curves (cf. Section 7.2 for a discussion).
@@
As in conventional public-key infrastructures (PKIs), one typically requires a means to invalidate already issued credentials before their expiration. However, when desiring unlinkability this cannot be a simple serial number that is always checked and consequently revocation in such a scenario is more complex and adds additional overhead. While this overhead is typically manageable from a computational perspective [DBLP:conf/cms/LaponKDN11] it might introduce additional issues and needs to be thoroughly evaluated. Loosely speaking, it typically mirrors the concept of a credential revocation list, but in a privacy-preserving way. One trivial approach that avoids such mechanisms and is conceptually simple is frequent re-issuing of credentials to users. This however adds a significant overhead on the issuer and puts some online requirements on the user. Moreover, this is dangerous from a privacy perspective, as a time-based correlation between issuing directly before using can drastically reduce the size of the anonymity set.
@@
Devices storing credentials can be lost or stolen. So it is important to make credentials recoverable or robust to device loss and unavailability respectively. One strategy is to back up credentials in a way so that they can be recovered when the device is not available anymore [DBLP:conf/acns/BaldimtsiCHKLN15]. An alternative approach is to consider sharing the credential among more than one device (in a threshold way) and consider threshold showing protocols, i.e., $t$ out of $n$ devices need to be availale, for credentials [cryptoeprint:2024/1874]. To the best of our knowledge this issue is rarely discussed, but we believe that it is an important that needs to be considered for practical deployment.
@@
A user might require to share access to resources and services with another person or among their different electronic devices. In practice, credentials are usually issued in a hierarchical manner, e.g., in a PKI there is a chain of certificates between the user certificate and a trusted root authority. In case this chain of issuers (or delegators) reveals sensitive information about the issuer’s organizational structure or the credential holder, one can rely on delegatable anonymous credentials [DBLP:conf/crypto/BelenkiyCCKLS09] (cf. [DBLP:journals/popets/MirSBM23] for a recent discussion on practically efficient schemes).
@@
In some use-cases it is required that while users are not identifiable, their action can still be linked together (even in case one uses multi-use credentials). Whenever this is required, it is possible to use context-specific pseudonyms. Here, from a unique value encoded in the credential (e.g., from its secret key) and a given context (string), a user always deterministically derives the same pseudorandom identifier that can be used as a pseudonym in this context. For instance, all actions of a user in the health domain are linkable, but are unlinkable to other domains.
@@
In a setting where a credential could have been obtained from multiple issuers, e.g., one country within the EU, it might be privacy critical to reveal the concrete issuer. The same holds when showing multiple credentials from several issuers, where the combination of issuers might already leak a lot of information. Consequently, the issuer-hiding concept [DBLP:conf/cans/BobolzEKRS21, DBLP:conf/pkc/Connolly0P22] allows to not reveal the issuer of a credential to the verifier during showing. Rather than revealing the issuer, the user just shows that an “issuer-policy”, i.e., membership in a set of issuers, defined by the verifier is satisfied.
@@
An issuer might be required not to learn certain attributes when issuing a credential. For instance, the user might want to include a secret key as an attribute into the credential, but the issuer is only allowed to get to know the public key. Moreover, it might be the case that the user wants to transfer a hidden attribute from one credential to another newly issued one without revealing anything about it (not even a function as in case of the secret and public key) to the issuer.
@@
Certain applications might require to make anonymity conditional and escrow identifying information with the showing of a credential (e.g., by encrypting it for a certain party). Consequently, this information can be opened by a third party when required and thus enabling re-identification of the otherwise anonymous user.
@@ 4.3 Construction Principles
There are different ways to construct anonymous credentials on a technical level. On a high-level, the user obtains a signature from the issuer on a set of attributes, which typically includes some secret of the user that is not revealed to the issuer. For showing, the user then demonstrates possession of such a valid signature from the issuer on a set of attributes that satisfies a certain relation. This can be just revealing a subset of the attributes or be some more complex relation. The important point is that the showing does not reveal the original signature from the issuer directly and thus cannot be linked to the issuing. Also, what is revealed during the showings is unlinkable.
Zero-knowledge proofs. If one wants to demonstrate a more complex relation, this typically requires the use of zero-knowledge proofs (ZKPs). We recall that ZKPs [DBLP:journals/siamcomp/GoldwasserMR89] allow one party (the prover) by interacting with another party (the verifier) to convince the latter that a statement (from some NP language) is true without revealing any additional information (the zero-knowledge property). At the same time, the prover is not able to make the verifier accept proofs about false statements (the soundness property). Often it is important to remove interaction in that the prover only needs to compute a single message (a proof), which can then be verified by everyone. These are so called non-interactive zero-knowledge (NIZK) proofs [DBLP:conf/stoc/BlumFM88].
Dedicated constructions. Anonymous credentials with a focus on obtaining highly efficient instantiations can be constructed from specific types of signatures schemes. They can be based on signatures with efficient protocols like CL [DBLP:conf/crypto/CamenischL04], BBS [DBLP:conf/crypto/BonehBS04, DBLP:conf/eurocrypt/TessaroZ23a], BBS+ [DBLP:conf/scn/AuSM06, DBLP:conf/trust/CamenischDL16] or PS [DBLP:conf/ctrsa/PointchevalS16] signatures, which are specific signature schemes that are compatible with commitment schemes and efficient zero-knowledge proofs. Alternatively, they can be based on signatures with specific randomization properties which allow to avoid zero-knowledge proofs [DBLP:conf/asiacrypt/Verheul01, DBLP:journals/joc/FuchsbauerHS19], or on specific redactable signatures that allow to redact signed messages [DBLP:conf/asiacrypt/CamenischDHK15, DBLP:conf/pkc/Sanders20].
@@
Generic constructions. A second and more generic way is to construct them from any signature scheme, e.g., standardized and widely deployed ones like the Elliptic Curve Digital Signature Algorithm (ECDSA), and a non-interactive zero-knowledge (NIZK) proof system. While generically this does not yield concretely efficient (and actually rather theoretic) constructions, in the recent decade there has been an enormous progress in the field of NIZK proofs and in particular on zero-knowledge succinct non-interactive arguments of knowledge (zk-SNARKs) (cf. [Nitulescu2020zkSNARKsAG, SEC-030]). Latter are succinct NIZK proofs, where loosely speaking succinctness means that the proofs are short and the verification of proofs is efficient. Using recent advances in this directions makes the construction of such anonymous credential schemes relying on existing signatures such as ECDSA practically feasible. There are very recent works by Google [cryptoeprint:2024/2010] and Microsoft [cryptoeprint:2024/2013] leveraging zk-SNARKs to turn existing credentials (i.e., ECDSA and ECDSA or RSA signatures respectively) into anonymous credentials.

### v2

@@ 4 Anonymous Credentials: The Theory
We now first want to clarify what we mean by “privacy-preserving authentication”. Essentially, this is a signature-based authentication approach that supports $i)$ selective disclosure and $ii)$ unlinkability in all the flavors that we have discussed. Consequently, even if IdP and RPs collude, it is not possible to track users. Note that this makes the user decide which information (e.g., attributes) are revealed and provide a strong unlinkability guarantee by default. As already mentioned in Section 1, there are different types of anonymous credentials with a set of different features as well as (closely) related primitives (cf. [DBLP:conf/secsr/KakviMPQ23, cryptoeprint:2023/1039]). In the following we will consider a simple abstract concept and will then discuss different (basic) features in Section 4.3.
@@ 4.1 The Abstract Concept
IssuerUserVerifier 1Verifier 2Issuing (attributes $A$)Presentation of $A_{1}\subseteq A$Presentation of $A_{2}\subseteq A$///Presentations are unlinkable
@@
At the end of this interaction, the user has a credential, which can be seen as a signature of the issuer on attributes $A$. Then, the user can perform a presentation protocol with verifiers, where in each interaction a subset $A_{i}\subseteq A$ of the attributes in the credential can be revealed. For simplicity, here we use a simple selective disclosure of attributes, but this could actually be a more complex predicate over a subset of attributes (as discussed in Section 4.3). For instance, a user might want to consume some content online where they need to demonstrate that they satisfy some age rating or that they are residents of some country. Note that in those cases revealing the full identity of the user is typically not necessary. If a verifier accepts such a presentation protocol, then it can be sure it interacts with a user who has a valid credential from the issuer for the subset of attributes that were revealed during the presentation protocol. Note that as long as none of the attributes presented are identifying, and there is more than one user who could potentially also reveal the same set of attributes, the user stays anonymous. Note that in Figure 2 we can also have that both verifiers are the same entity.
@@
Security. What we request from an anonymous credential scheme is that it is unforgeable, that is, it is not possible for a user to demonstrate some set of attributes for which they did not obtain a credential. Moreover, the scheme is unlinkable, which means that it is not possible to link together the presentations of the same user. We note that for any candidate scheme considered to be deployed in practice it is imperative to come with a rigorous proof of security in a realistic security model.
@@
Extensions to this simple model. As briefly mentioned above, realistic scenarios often involve multiple issuers — a situation sometimes referred to as a multi-authority anonymous credential system. In such a setting, every issuer has an independent signing key to issue credentials, and a presentation might then involve more than one credential. Such an extension is typically straightforward, but if one targets at specific properties, e.g., constant-size credential presentations irrespective of the number of issuers, dedicated constructions exist [DBLP:conf/scn/HebantP22, DBLP:conf/ccs/MirBGLS23].
@@
Another flavor of having multiple issuers is that of threshold issuance anonymous credentials [DBLP:conf/ndss/SonninoABMD19, DBLP:conf/sp/DoernerKLST23]. Here, rather than every issuer holding an independent signing key, all the credentials are issued under a single signing key. However, the key is shared among all the issuers using a $(k,n)$-threshold secret sharing technique. In such a setting, there are $n$ issuers each holding a share of the secret key and $k$ issuers need to collaborate to be able to issue a credential. Note that this is a means to distribute the trust in the issuers and adds security, as it is required that even leaking $t<k$ secret key shares does not allow one to learn the signing key. Moreover, it adds robustness as the availability of any $k$ out of the $n$ issuers is sufficient to issue credentials.
4.2 Anonymous Credentials and Distributed Ledger Technologies
We note that today there is a trend to move away from a centralized setting towards a decentralized identity. A popular concept in this decentralized identity space is that of self-sovereign identity (SSI), with Sovrin [^7] https://sovrin.org being a prominent example that leverages distributed ledger technologies (DLTs) and concepts from this domain such as decentralized identifiers (DIDs).
When talking about the use of concepts from DLTs together with anonymous credentials, there are two different approaches that should not be confused. First, proposals such as decentralized anonymous credentials [DBLP:conf/ndss/Garman0M14] or zk-creds [DBLP:conf/sp/RosenbergWGM23], that crucially use DLTs to realize novel types of anonymous credentials. This is done with the goal of not requiring explicit credential issuers [DBLP:conf/ndss/Garman0M14], or they are designed to solely rely on existing identity documents and thus do not require issuers to issue additional signature-based credentials [DBLP:conf/sp/RosenbergWGM23].
Second, the practical implementation of existing anonymous credentials using technologies provided by DLTs such as DIDs. For instance, in the context of anonymous credentials, DIDs commonly provide a registry-based mechanism to “anchor” the identities of credential issuers, holders, and verifiers. Anonymous credential solutions might also leverage DLT to store credential schemas and registries for credential revocation. Consequently, even though DLTs and DIDs can function independently of anonymous credentials (and vice versa), they are often tightly intertwined in practice, and many standardization initiatives have evolved in parallel across both areas. Examples are Hyperledger AnonCreds developed alongside Hyperledger Indy, and the World Wide Web Consortium (W3C) Verifiable Credentials standards developed in tandem with DID standards (cf. Section 5 for more details).
Moreover, the European Union Digital Identity Wallet (EUDIW) later discussed in Section 6, although not related to DLTs and in contrast to Sovrin not yet related to anonymous credentials, can be viewed as an approach towards SSI. At a high level, users are collecting certified attributes, i.e., verifiable credentials, from different sources and then presenting (subsets of) verifiable credentials from this collection.
@@ 4.3 Features of Anonymous Credentials
In the following, we present some basic, distinguishing features of anonymous credentials, as well as a list of extended features. We try to be comprehensive in our presentation, but do not claim to be exhaustive. See also [DBLP:journals/istr/CamenischDELNPP14] for a discussion of features.
@@
In the single-use case, if a credential is used more than once, different presentations become linkable. Multi-use credentials support an unlimited number of unlinkable presentations. When talking about anonymous credentials, one typically implicitly assumes multi-use ones.
@@
Typically it should be possible to encode attributes into an anonymous credential. However, over the recent years some use cases appeared in industry in which attributes are not really required. The latter concept is then typically called anonymous tokens, starting with Privacy Pass [DBLP:journals/popets/DavidsonGSTV18], which are built from blind signatures or related primitives such as oblivious pseudo-random functions (OPRFs).
@@
An issue with any digital credential is that it can easily be copied and thus illegally distributed. This gets particularly delicate in case of privacy-preserving authentication mechanisms where the identity will not be revealed. While there is no panacea, a typical approach is to bind it to hardware [DBLP:conf/crypto/Brands93, DBLP:conf/ccs/HanzlikS21], as done within the related primitive of direct anonymous attestation (DAA) [DBLP:conf/ccs/BrickellCC04] built within every Trusted Platform Module (TPM), available in most modern computers nowadays. An alternative approach is to force users to share a valuable secret whenever they share their credentials [DBLP:conf/eurocrypt/CamenischL01], a concept which, however, seems rather hard to implement in practice. We stress that even when credentials are bound to hardware and thus a device, this does not yet prevent the sharing of the device, e.g., lending it to a family member or friend. There are recent approaches to involve biometrics to prevent this problem. Hesse et al. suggest to use a dedicated hardware token that displays a photo of the credential holder and which is bound to an anonymous credential [DBLP:conf/uss/HesseSS23]. Adams suggested to directly integrate a biometric template as attribute into the anonymous credential [DBLP:journals/scn/Adams11]. While this approach has recently been shown to be practical [DBLP:journals/compsec/RodriguezKS24], it is best suited for physical access control (or requires additional trust assumptions).
Expressiveness of attribute presentations.
Credentials may either only allow to reveal or withhold attributes (this can be thought of redacting the non-presented attributes) or be able to prove arbitrary statements about attributes encoded in the credential, e.g., attribute country represents any of the EU countries and attribute birthdate is so that the holder is above 18 years old. The latter requires the use of zero-knowledge proofs (cf. Section 4.4), while from a construction perspective former can be realized without them.
@@
Typically, credentials are publicly verifiable using the public key of the issuer. However, there are also constructions [DBLP:conf/ccs/ChaseMZ14, DBLP:journals/iacr/Orru24] and use cases, e.g., private groups in the Signal messenger [DBLP:conf/ccs/ChasePZ20], where the issuer is identical to the verifier and verification uses the secret of the issuer. This can allow for more efficient constructions that can avoid the use of bilinear groups and thus specific pairing-friendly elliptic curves (cf. Section 7.2 for a discussion).
@@
As in conventional public-key infrastructures (PKIs), one typically requires a means to invalidate already issued credentials before their expiration. However, when desiring unlinkability this cannot be a simple serial number that is always checked and consequently revocation in such a scenario is more complex and adds additional overhead. While this overhead is typically manageable from a computational perspective [DBLP:conf/cms/LaponKDN11] it might introduce additional issues and needs to be thoroughly evaluated. Loosely speaking, it typically mirrors the concept of a credential revocation list, but in a privacy-preserving way. One trivial approach that avoids such mechanisms and is conceptually simple is frequent re-issuing of credentials to users. This, however, adds a significant overhead on the issuer and puts some online requirements on the user. Moreover, this can be dangerous from a privacy perspective, as a time-based correlation between issuing directly before using can drastically reduce the size of the anonymity set.
@@
Devices storing credentials can be lost or stolen. So it is important to make credentials recoverable or robust to device loss and unavailability respectively. One strategy is to back up credentials in a way so that they can be recovered when the device is not available anymore [DBLP:conf/acns/BaldimtsiCHKLN15]. An alternative approach is to consider sharing the credential among more than one device (in a threshold way) and consider threshold presentation protocols for credentials [cryptoeprint:2024/1874], i.e., $t$ out of $n$ devices need to be available. To the best of our knowledge this issue is rarely discussed, but we believe that it is an important one that needs to be considered for practical deployment.
@@
A user might require to share access to resources and services with another person or among their different electronic devices. In practice, credentials are usually issued in a hierarchical manner, e.g., in a PKI there is a chain of certificates between the user certificate and a trusted root certification authority. In case this chain of issuers (or delegators) reveals sensitive information about the issuer’s organizational structure or the credential holder, one can rely on delegatable anonymous credentials [DBLP:conf/crypto/BelenkiyCCKLS09] (cf. [DBLP:journals/popets/MirSBM23] for a recent discussion on practically efficient schemes).
@@
In some use cases it is required that while users are not identifiable, their action can still be linked together (even in case of multi-use credentials). Whenever this is required, it is possible to use context-specific pseudonyms. Here, from a unique value encoded in the credential (e.g., from its secret key) and a given context (string), a user always deterministically derives the same pseudorandom identifier that can be used as a pseudonym in this context. For instance, all actions of a user in the health domain are linkable, but are unlinkable to other domains.
@@
In a setting where a credential could have been obtained from multiple issuers, e.g., one country within the EU, it might be privacy critical to reveal the concrete issuer. The same holds when presenting multiple credentials from several issuers, where the combination of issuers might already leak a lot of information. Consequently, the issuer-hiding concept [DBLP:conf/cans/BobolzEKRS21, DBLP:conf/pkc/Connolly0P22] allows to not reveal the issuer of a credential to the verifier during presentation. Rather than revealing the issuer, the user just demonstrates that an “issuer-policy”, i.e., membership in a set of issuers, which is defined by the verifier, is satisfied.
@@
An issuer might be required not to learn certain attributes when issuing a credential. For instance, users typically need to include a secret key as an attribute into the credential, but the issuer is only allowed to get to know the public key. Moreover, it might be the case that the user wants to transfer a hidden attribute from one credential to another newly issued one without revealing anything about it (not even a function as in case of the secret and public key) to the issuer.
@@
Certain applications might require to make anonymity conditional and to escrow identifying information with the presentation of a credential (e.g., by encrypting it for a certain party). Consequently, this information can be opened by a third party when required and thus enables re-identification of the otherwise anonymous user.
@@ 4.4 Construction Principles
There are different ways to construct anonymous credentials on a technical level. At a high level, the user obtains a signature from the issuer on a set of attributes, which typically includes some secret of the user that is not revealed to the issuer. For presentation, the user then demonstrates possession of such a valid signature from the issuer on a set of attributes that satisfies a certain relation. This can be just revealing a subset of the attributes or be some more complex relation. The important point is that the presentation does not reveal the original signature from the issuer directly and thus cannot be linked to the issuing. Also, what is revealed during the presentations is unlinkable.
Zero-knowledge proofs. If one wants to demonstrate a more complex relation, this typically requires the use of zero-knowledge proofs (ZKPs). We recall that ZKPs [DBLP:journals/siamcomp/GoldwasserMR89] allow one party (the prover) by interacting with another party (the verifier) to convince the latter that a statement (from some NP language) is true without revealing any additional information (the zero-knowledge property). At the same time, the prover is not able to make the verifier accept proofs about false statements (the soundness property). Often it is important to remove interaction in that the prover only needs to compute a single message (a proof), which can then be verified by everyone. These are so-called non-interactive zero-knowledge (NIZK) proofs [DBLP:conf/stoc/BlumFM88]. It should be noted that deploying ZKPs or NIZKs introduces additional computational and implementation complexities and requires a certain level of cryptographic know-how.
Dedicated constructions. Anonymous credentials with a focus on obtaining highly efficient instantiations can be constructed from specific types of signatures schemes. They can be based on signatures with efficient protocols like CL [DBLP:conf/crypto/CamenischL04], BBS [DBLP:conf/crypto/BonehBS04, DBLP:conf/eurocrypt/TessaroZ23a], BBS+ [DBLP:conf/scn/AuSM06, DBLP:conf/trust/CamenischDL16] or PS [DBLP:conf/ctrsa/PointchevalS16] signatures, which are specific signature schemes that are compatible with commitment schemes and efficient ZKPs. Alternatively, they can be based on signatures with specific randomization properties which allow to avoid ZKPs [DBLP:conf/asiacrypt/Verheul01, DBLP:journals/joc/FuchsbauerHS19], or on specific redactable signatures that allow to redact signed messages [DBLP:conf/asiacrypt/CamenischDHK15, DBLP:conf/pkc/Sanders20].
@@
Generic constructions. A second and more generic way is to construct them from any signature scheme, e.g., standardized and widely deployed ones like the Elliptic Curve Digital Signature Algorithm (ECDSA), and a NIZK proof system. While generically this does not yield concretely efficient (and actually rather theoretic) constructions, in the recent decade there has been enormous progress in the field of NIZK proofs and in particular on zero-knowledge succinct non-interactive arguments of knowledge (zk-SNARKs) (cf. [Nitulescu2020zkSNARKsAG, SEC-030]). zk-SNARKs are succinct NIZK proofs, where loosely speaking succinctness means that the proofs are short and the verification of proofs is efficient. Using recent advances in this directions makes the construction of such anonymous credential schemes relying on existing signatures such as ECDSA practically feasible. There are very recent works by Google [cryptoeprint:2024/2010] and Microsoft [cryptoeprint:2024/2013] leveraging zk-SNARKs to turn existing credentials (i.e., ECDSA and ECDSA or RSA signatures respectively) into anonymous credentials.